OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

TheSmartOne (thesmartone):

Ok, so which equation do you want to handle first? :)

TheSmartOne (thesmartone):

Also, welcome to OpenStudy! :D @casey321123 @~

OpenStudy (anonymous):

i am not really understanding how to do this problem

TheSmartOne (thesmartone):

ok, let's take the #1 equation :) which is 2x + y = 1

TheSmartOne (thesmartone):

now let's solve for y in that equation :D 2x + y = 1 Subtract 2x on both sides, what do you get?

OpenStudy (anonymous):

y=1-2x

TheSmartOne (thesmartone):

correct, now we plug that equation into the second equation

TheSmartOne (thesmartone):

4x + 2y = −2 so in place of y over there, we plut in 1-2x so we get: 4x + 2(1-2x) = -2 solve for x :)

OpenStudy (anonymous):

x=-2

TheSmartOne (thesmartone):

4x + 2(1-2x) = -2 First use the distributive property a(b-c) = ab - ac so 2(1-2x) = ?

TheSmartOne (thesmartone):

not sure how you got x = -2

TheSmartOne (thesmartone):

so let's do it step by step :)

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

so what is 2(1-2x) = ?

OpenStudy (anonymous):

idk what to do

TheSmartOne (thesmartone):

Use the distributive formula a(b-c) = ab - ac so 2(1-2x) = ?? What is 2 x 1 = ? what is 2 x 2 = ?

OpenStudy (anonymous):

2 and 4

TheSmartOne (thesmartone):

ok, good so 2(1-2x) = 2 - 4x

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

4x + 2 - 4x = -2 What is 4x - 4x = ?

OpenStudy (anonymous):

x

TheSmartOne (thesmartone):

nope 4x - 4x = (x)(4-4) What is 4-4 =?

OpenStudy (anonymous):

0

TheSmartOne (thesmartone):

correct so 4x - 4x = (4-4)(x) = 0x = 0

OpenStudy (anonymous):

no solution

TheSmartOne (thesmartone):

so tell me, @casey321123 does \(\sf -2 = 2\)?

TheSmartOne (thesmartone):

exactly! It's no solution! :D

TheSmartOne (thesmartone):

when we get no solution, that means, that the two lines are parallel. :) Great job! :D